OpenStudy (miguel980):

Which two integers have a difference of 24 and a quotient of -7?

OpenStudy (anonymous):

So, you have two variables. For simplicity's sake, let's name then x and y. Condition 1) They have a difference of 24. \[x - y = 24\] Condition 2) They have a quotient of -7 \[\frac{ x }{ y } = -7\] From here, there are multiple algebraic options. But for now, let's just use simple substitution. \[x = -7y\] \[(-7y) - y = 24\] \[-8y = 24\] \[y = -\frac{ 24 }{ 8 } = ~?\] How would you further simplify this value? @Miguel980

OpenStudy (anonymous):

After you find the y-value, substitute it into either of the two equations above to find your x-value. By then, you'll have both your unknown integers.